Jim Ryan, an owner of a Burger King restaurant, assumes that his restaurant will need a new roof in 5 years. He estimates the roof will cost him $10,300 at that time.


What amount should Jim invest today at 12% compounded quarterly to be able to pay for the roof? (Do not round intermediate calculations. Round your answer to the nearest cent.)
Answer #1

Answer

quarterly effective interest rate =annual nominal rate /number of compounding =12/4=3%=i

n=number of quarters =5*4=20

FV=future value=10300

PV=presentt value =?

PV=FV*(1+i)^(-n)

=10300*(1.03^-20)

=5702.86027

=5702.86

the value is $5702.86